# P3-LLM: An Integrated NPU-PIM Accelerator for LLM Inference Using Hybrid Numerical Formats

### v2

@@ IV-D Activation Quantization
Similar to the key cache, activations are also known to exhibit outlier channels [llm-int8], which makes standard integer quantization error-prone. To reduce the quantization error of integer activations, prior works mainly leverage two techniques to suppress outliers. The first is Hadamard transformation, which applies a Hadamard rotation matrix to the input activation before performing integer quantization [quarot, qserve]. However, the online Hadamard transformation can bring considerable runtime overhead due to additional memory access and matrix operations [ecco]. The second is smoothing, which migrates the quantization difficulty of outlier activation channels to the corresponding weight channels [smoothquant]. Unfortunately, the migration of quantization difficulty can significantly increase the error of low-precision weights.
To address these issues, we propose leveraging an 8-bit numerical format that remains well-suited for activations even in the presence of outliers. Interestingly, we find that the standard FP8-E4M3 format can achieve near-lossless accuracy for per-token activation quantization, owing to its wide numerical range. Table III shows the perplexity of activation quantization using FP8-E4M3 and SmoothQuant-based [smoothquant] INT8. SmoothQuant causes large perplexity loss under 4-bit weights. This is expected, because weights are already difficult to quantize at 4 bits, and activation smoothing migrates additional quantization difficulty to weights. On the other hand, FP8-E4M3 significantly outperforms SmoothQuant as its large numerical range can accommodate outliers without affecting weight quantization.
@@ V P3-LLM Architecture
This section outlines the architectural details of P3-LLM, which features: (a) a high-throughput and area-efficient PCU co-designed to support hybrid numerical formats, (b) an efficient operator mapping that maps most memory-intensive operators to the low-precision PIM hardware during decoding, (c) a quantized matrix multiplication dataflow that applies operator fusion to minimize the runtime overhead of quantization and dequantization.
@@ V-C Optimization on Quantized Computation Flow
Quantized Computation Flow. As shown in Fig. 6(c), P3-LLM implements operator fusion to minimize the runtime overhead of (de)quantization, which needs to be executed on NPU using high-precision arithmetic. For linear layers, the dequantization scaling is performed after matrix multiplication. For $Q\cdot K^{\text{T}}$}, since the post-RoPE key cache contains per-channel smoothing factors ($S_{\text{SF}}$), we fuse $S_{\text{SF}}$ into query via element-wise multiplication prior to FP8 quantization. Similarly, for $P\cdot V$}, we fuse the per-value-head scaling factors ($S_{V}$) into attention-scores before FP8 quantization. One potential issue of fusing $S_{V}$ is that it may cause the resulting attention-scores’ magnitude to become larger than one, which is outside the range of FP8-S0E4M4. To prevent this, we apply a second-level scaling that divides $S_{V}$ with its maximum, resulting in $S_{V}\in[0,1]$}. Note that the quantization scaling factor is always positive by definition in Eq. 1, so fusing $S_{V}$ still produces unsigned attention-scores. This second-level scaling factor is later multiplied back into the result of $P\cdot V$}.
@@
    For weight-activation quantization, QuaRot and QServe maintain query and attention-scores in FP16. Whereas P3-LLM quantizes attention-scores to FP8-S0E4M4 across all models, and quantizes query to FP8-E4M3 for Llama-3 and Mistral.
@@ VI-A Experimental Methodology
Algorithm Baselines. We compare P3-LLM with SoTA quantization algorithms, including Oaken [oaken], QuaRot [quarot], and QServe [qserve]. Oaken is a 4-bit KV-cache quantization method that determines the key-cache outlier threshold offline using a calibration dataset. We compare Oaken with P3-LLM under KV-cache-only quantization for a fair comparison. QuaRot and QServe implement aggressive W4A8KV4 quantization. For all baseline algorithms, we use their official GitHub repository to reproduce the accuracy.
@@
TABLE V: Accuracy ($\uparrow$) of reasoning tasks under different quantization methods. Oaken and the top P3-LLM row use KV4 quantization. QuaRot and QServe use W4A8KV4 quantization, while the bottom P3-LLM row uses W4A8KV4P8 quantization.
@@ VI-B Model Performance
Figure 8: Normalized layer-wise key-cache quantization error of Llama-2-7B on Wikitext-2 and C4 datasets. Oaken and QServe use Wikitext-2 and Pile as calibration datasets, respectively.
@@
Analysis of Key-cache Quantization Error. Fig. 8 presents the normalized key-cache quantization error of Oaken, QServe, and P3-LLM on the Wikitext-2 and C4 datasets across all layers of Llama-2-7B. For each layer, we normalize the error to the average magnitude of all samples in the dataset. Oaken and QServe rely on Wikitext-2 and Pile [pile] as calibration datasets to address the key-cache outliers. Although Oaken has similar quantization error as P3-LLM on Wikitext-2, its error becomes more pronounced on C4 due to overfitting. QServe has the highest error since the calibration statistics obtained from Pile does not generalize well to both Wikitext-2 and C4. In contrast, the dynamic input-aware smoothing of P3-LLM yields the lowest quantization error, as it does not overfit to any calibration dataset.
@@ VII Related Works
Custom Quantization Formats. Recently, there has been significant research on custom numerical formats for quantization in the architecture community [microscaling, fp8, nvfp4, olive, ant, anda, mant, bitmod, fp8_custom]. OliVe [olive] proposes to quantize LLM outliers using an adaptive biased-float format. ANT [ant] introduces a new data type that better adapts to the per-tensor value distribution of different model layers. MANT [mant] further extends ANT with a mathematically adaptive data type for group-wise quantization. Although these proposals achieve good hardware efficiency, they struggle to outperform existing algorithmic solutions under the aggressive W4A8KV4 configuration [qserve, quarot]. On the contrary, P3-LLM introduces a novel operand-dependent quantization scheme with hybrid numerical formats, achieving higher accuracy than SoTA algorithmic solutions. In addition to academic proposals, custom quantization formats have been widely adopted by industry. The microscaling format [microscaling], standardized by Open Compute Project, applies a shared 8-bit exponent to a group of low-precision operands. Both NVIDIA and AMD add support for FP4 and FP8 formats in their flagship GPUs [amd-gpu, blackwell-gpu].
PIM-based LLM Inference Systems. There is an abundance of works on LLM inference using NPU-PIM heterogeneous accelerators [neupim, pimba, ianus, duplex, plain, lp-spec, samsung-aim, sk-aim, attacc]. Samsung and SK Hynix have commercialized PIM products based on DDR and HBM technology [samsung-aim, sk-aim]. AttAcc [attacc], IANUS [ianus], and NeuPIMs [neupim] explore operator mapping and scheduling to improve the utilization of NPU and PIM during inference. These works focus exclusively on FP16 LLM inference, making them orthogonal to our quantization proposal. On the other hand, Pimba [pimba], PLAIN [plain], and LP-Spec [lp-spec] design low-precision arithmetic units to enhance the throughput of PIM. However, they mainly target conservative quantization configurations such as W8A8 and W6A6. In contrast, our work explores the more aggressive W4A8KV4P8 quantization to achieve an optimal trade-off between model accuracy, memory footprint, and computational efficiency.
@@ VIII Conclusion
In this paper, we present P3-LLM, an algorithm-hardware co-design framework for efficient LLM inference on heterogeneous NPU-PIM systems. On the algorithm side, P3-LLM develops a novel mixed-precision quantization scheme with hybrid numerical formats to quantize all major LLM operands, while maintaining superior model accuracy. On the hardware side, P3-LLM is equipped with a high-throughput and area-efficient PCU architecture, which contains lightweight mixed-precision processing elements optimized for the proposed quantization scheme. Our evaluation demonstrates that P3-LLM significantly outperforms existing KV-cache-only and weight-activation quantization methods. Compared to SoTA LLM accelerators HBM-PIM, Ecco, and Pimba, P3-LLM yields $4.9\times$, $2.0\times$, and $3.4\times$ speedups, respectively, while offering an optimal trade-off between model accuracy and hardware efficiency.

### v3

@@ IV-D Activation Quantization
Similar to the key cache, activations are also known to exhibit outlier channels [llm-int8], which makes standard integer quantization error-prone. To reduce the quantization error of integer activations, prior works mainly leverage two techniques to suppress outliers. The first is Hadamard transformation, which applies a Hadamard rotation matrix to activations before performing quantization [quarot]. However, the online Hadamard transformation can bring considerable runtime overhead due to additional memory access and matrix operations [ecco]. The second is smoothing, which migrates the quantization difficulty of outlier activation channels to the corresponding weight channels [smoothquant]. Unfortunately, the migration of quantization difficulty can significantly increase the error of low-precision weights.
To address these issues, we propose leveraging an 8-bit format that remains well-suited for activations even in the presence of outliers. Inspired by prior studies [zeroq_fp, fp8_vs_int8], we explore FP8 as an alternative format for activation quantization without Hadamard transformation and smoothing. We find that the standard FP8-E4M3 format can achieve near-lossless accuracy for quantizing activations, owing to its wide numerical range. Table III shows the perplexity of per-token activation quantization using FP8-E4M3 and SmoothQuant-based [smoothquant] INT8. SmoothQuant causes large perplexity loss under 4-bit weights. This is expected, because weights are already difficult to quantize at 4 bits, yet SmoothQuant migrates additional difficulty to weights. On the other hand, FP8-E4M3 significantly outperforms SmoothQuant as its large numerical range can accommodate outliers without affecting weight quantization.
@@ V P3-LLM Architecture
This section outlines the architectural details of P3-LLM, which features: (a) a high-throughput and area-efficient PCU co-designed to support hybrid numerical formats, (b) an efficient operator mapping that maps memory-intensive operators to the low-precision PIM hardware during decoding, (c) a quantized matrix multiplication flow that applies operator fusion to minimize the runtime overhead of dequantization.
@@ V-C Optimization on Quantized Computation Flow
Quantized Computation Flow. As shown in Fig. 6(c), P3-LLM performs quantized matrix multiplication on the low-precision PCU, and implements operator fusion to minimize the runtime overhead of dequantization, which needs to be executed on NPU using high-precision arithmetic. For linear layers, the dequantization scaling is performed after matrix multiplication. For $Q\cdot K^{\text{T}}$}, since the post-RoPE key cache contains per-channel smoothing factors ($S_{\text{SF}}$), we fuse $S_{\text{SF}}$ into query via element-wise multiplication prior to FP8 quantization. Similarly, for $P\cdot V$}, we fuse the per-value-head scaling factors ($S_{V}$) into attention-scores before FP8 quantization. One potential issue of fusing $S_{V}$ is that it may cause the resulting attention-scores’ magnitude to become larger than one, which is outside the range of FP8-S0E4M4. To prevent this, we apply a second-level scaling that divides $S_{V}$ with its maximum $S_{V\text{max}}$, resulting in $S_{V}\in[0,1]$}. Note that the quantization scaling factor is always positive by definition in Eq. 1, so fusing $S_{V}$ still produces unsigned attention-scores. This second-level scaling factor $S_{V\text{max}}$ is later multiplied back into the result of $P\cdot V$}.
@@
    For weight-activation quantization, QuaRot and QoQ maintain query and attention-scores in FP16. Whereas P3-LLM quantizes attention-scores to FP8-S0E4M4 across all models, and quantizes query to FP8-E4M3 for Llama-3 and Mistral.
@@ VI-A Experimental Methodology
Algorithm Baselines. We compare P3-LLM with SoTA quantization algorithms, including Oaken [oaken], QuaRot [quarot], and QoQ [qserve]. Oaken is a 4-bit KV-cache quantization method that determines the key-cache outlier threshold offline using a calibration dataset. We compare Oaken with P3-LLM under KV-cache-only quantization for a fair comparison. QuaRot and QoQ implement aggressive W4A8KV4 quantization. For all baseline algorithms, we use their official GitHub repository to reproduce the accuracy.
@@
TABLE V: Accuracy ($\uparrow$) of reasoning tasks under different quantization methods. Oaken and the top P3-LLM row use KV4 quantization. QuaRot and QoQ use W4A8KV4 quantization, while the bottom P3-LLM row uses W4A8KV4P8 quantization.
@@ VI-B Model Performance
Figure 8: Normalized layer-wise key-cache quantization error of Llama-2-7B on Wikitext-2 and C4 datasets. Oaken and QoQ use Wikitext-2 and Pile as calibration datasets, respectively.
@@
Analysis of Key-cache Quantization Error. Fig. 8 presents the normalized key-cache quantization error of Oaken, QoQ, and P3-LLM on the Wikitext-2 and C4 datasets across all layers of Llama-2-7B. For each layer, we normalize the error to the average magnitude of all samples in the dataset. Oaken and QoQ rely on Wikitext-2 and Pile [pile] as calibration datasets to address the key-cache outliers. Although Oaken has similar quantization error as P3-LLM on Wikitext-2, its error becomes more pronounced on C4 due to overfitting. QoQ has the highest error since the calibration statistics obtained from Pile does not generalize well to both Wikitext-2 and C4. In contrast, the dynamic input-aware smoothing of P3-LLM yields the lowest quantization error, as it does not overfit to any calibration dataset.
@@ VII Related Works
Custom Quantization Formats. Recently, there has been significant research on custom numerical formats for quantization in the architecture community [mant, bitmod, blockdialect, ant, fp8, microscaling, olive, mx_plus, spark, amove, microscopiq, anda]. OliVe [olive] proposes to quantize LLM outliers using an adaptive biased-float format. ANT [ant] introduces a new data type that better adapts to the per-tensor value distribution of different model layers. MANT [mant] further extends ANT with a mathematically adaptive data type for group-wise quantization. Although these proposals achieve good hardware efficiency, they struggle to outperform existing algorithmic solutions under the aggressive W4A8KV4 configuration [qserve, quarot]. On the contrary, P3-LLM introduces a novel operand-dependent quantization scheme using hybrid numerical formats, achieving higher accuracy than SoTA algorithmic solutions. In addition to academic proposals, custom quantization formats have been widely adopted by industry. The microscaling format [microscaling], standardized by Open Compute Project, applies a shared 8-bit exponent to a group of low-precision operands. Both NVIDIA and AMD add support for FP4 and FP8 formats in their flagship GPUs [amd-gpu, blackwell-gpu]. The latest NVIDIA Blackwell architecture introduces the NVFP4 format [nvfp4], which enhances the standard FP4 data type with a two-level scaling strategy.
PIM-based LLM Inference Systems. There is an abundance of works on LLM inference using NPU-PIM heterogeneous accelerators [neupim, pimba, ianus, duplex, plain, lp-spec, samsung-aim, sk-aim, attacc, orches]. Samsung and SK Hynix have commercialized PIM products based on DDR and HBM technology [samsung-aim, sk-aim]. AttAcc [attacc], IANUS [ianus], and NeuPIMs [neupim] explore operator mapping and scheduling to improve the utilization of NPU and PIM during inference. These works focus exclusively on FP16 LLM inference, making them orthogonal to our quantization proposal. On the other hand, Pimba [pimba], PLAIN [plain], and LP-Spec [lp-spec] design low-precision arithmetic units to enhance the throughput of PIM. However, they mainly target conservative quantization configurations such as W8A8 and W6A6. In contrast, our work explores the more aggressive W4A8KV4P8 quantization to achieve an optimal trade-off between model accuracy, memory footprint, and computational efficiency.
@@ VIII Conclusion
In this paper, we present P3-LLM, an algorithm-hardware co-design framework for efficient LLM inference on heterogeneous NPU-PIM accelerators. On the algorithm side, P3-LLM develops a novel quantization scheme using hybrid numerical formats to quantize all major LLM operands, while maintaining superior model accuracy. On the hardware side, P3-LLM is equipped with optimized low-precision dataflow and an area-efficient PCU architecture to support the proposed quantization scheme. Our evaluation demonstrates that P3-LLM significantly outperforms existing KV-cache-only and weight-activation quantization methods. Compared to SoTA LLM accelerators HBM-PIM, Ecco, and Pimba, P3-LLM yields $4.9\times$, $2.0\times$, and $3.4\times$ speedups, respectively.